# Electron-phonon coupling in copper-substituted lead phosphate apatite

### v1

@@ Abstract
Recent reports of room-temperature, ambient pressure superconductivity in copper-substituted lead phosphate apatite, commonly referred to as LK99, have prompted numerous theoretical and experimental studies into its properties. As the electron-phonon interaction is a common mechanism for superconductivity, the electron-phonon coupling strength is an important quantity to compute for LK99. In this work, we compare the electron-phonon coupling strength among the proposed compositions of LK99. The results of our study are in alignment with the conclusion that LK99 is not a likely candidate for room-temperature superconductivity if electron-phonon interaction is to serve as the mechanism.
@@ I Introduction
A subsequent theoretical study computed the bulk band structure of LK99 considering two distinct locations for copper substitution, denoted Pb(1) and Pb(2), in the lead-phosphate apatiteGriffin (2023), Pb(PO)X, where X=O, OH. In each case, the existence of nearly-flat bands at the Fermi energy was demonstrated. The presence of such bands increases the density of states at the Fermi energy increasing the possibility of a mechanism for high-temperature superconductivity.
In the intervening months many questions about the nature of electronic states in LK99 were raised. Additional studies have been published with remarkable speed and detailCabezas-Escares et al. (2023); Hao and Fu (2024); Jiang et al. ; Lai et al. (2024); Korotin et al. (2023); Lee and Dai (2023); Shen et al. (2023), providing evidence that suggest LK99 is not a high-temperature superconductor. Among these studies, we are aware of only one work that has directly examined electron-phonon coupling strengthPaudyal et al. (2023). This is expected given the computational demands of such a quantity as well as the fact that many available density functional theory packages do not offer the capability for computation of electron-phonon coupling strength if a non-zero value of the Hubbard U is specified for a constituent atom in the compoundGiustino (2017).
In Ref. Paudyal et al. (2023), the electron-phonon coupling strength was computed from first-principles for PbCu(PO)O with Cu substituted at the Pb(2) site neglecting inclusion of a Hubbard U. It was shown that the electron-phonon coupling is weak. Even without inclusion of a Hubbard U for copper, this was a fascinating result as the electron-phonon interaction is one of the most common routes to the realization of superconductivity. It is thus a quantity of supreme importance and should be compared among the possible compositions of copper substituted lead phosphate apatite.
In this work we follow a recently established protocol to compute a proxy for the electron-phonon coupling strength, . This proxy involves computation of the electron-phonon coupling at the point only. In Ref. Sun et al. (2022), it is shown that,
where it is reasonable to estimate as a constant for a family of materials. It was further shown in Ref. Sun et al. (2022), that the magnitude of is generally sufficient for distinguishing systems likely to support superconductivity from those in which it is highly unlikely. As a result, we have computed for Pb(PO)X X=O, OH, with copper substitution at the Pb(1) site, allowing for a direct comparison between the compositions.
@@ II Results
Pb(PO)X belongs to space group and for X=O, OH respectively. Details of the structure before and after copper substitution can be found in Ref. Griffin (2023). The locations for copper substitution on the two in-equivalent Pb sites, labeled Pb(1) and Pb(2) are shown in Fig. 1. The relaxed structures for each of the compounds after copper substitution can be seen in Fig. LABEL:fig:Compounds. For further information regarding the crystal structure, bulk band structure, and more, we refer the reader to Refs. Griffin (2023); Cabezas-Escares et al. (2023); Shen et al. (2023); Jiang et al. .
The resulting phonon modes at the location are shown for PbCu(PO)X with Cu at the Pb(1) site for X=OH, O in Fig. LABEL:fig:Compounds. We immediately note that the absence of negative frequency modes for X=O, indicating that this compound is dynamically stable. In contrast, for X=OH negative phonon modes can be found. While this can indicate a dynamic instability, it is likely that such negative frequency modes would be removed through imposition of a stricter convergence threshold when relaxing the structure. This presents a computational challenge due to the lightness of the hydrogen atoms whose positions can be easily adjusted with minimal effects on the energy of the structure, slowing convergence.
The computed electron-phonon coupling strength at the location at the Fermi energy as a function of the smearing values utilized in the computation is shown in Fig. 2. In order to estimate for each compound, we utilize the value obtained in Ref. Paudyal et al. (2023) for PbCu(PO)O with Cu substituted at the Pb(2) site. We then determine , where
@@
In Ref. Paudyal et al. (2023), a superconducting critical temperature of was computed for PbCu(PO)O with Cu substituted at the Pb(2) site. Given the current results, a similar value should be obtained for substitution at the Pb(1) site.

### v2

@@ Abstract
Recent reports of room-temperature, ambient pressure superconductivity in copper-substituted lead phosphate apatite, commonly referred to as LK99, have prompted numerous theoretical and experimental studies into its properties. As the electron-phonon interaction is a common mechanism for superconductivity, the electron-phonon coupling strength is an important quantity to compute for LK99. In this work, we compare the electron-phonon coupling strength among the proposed compositions of LK99. The results of our study are in alignment with the conclusion that LK99 is a candidate for low-temperature, not room-temperature, superconductivity if electron-phonon interaction is to serve as the mechanism.
@@ I Introduction
A subsequent theoretical study computed the bulk band structure of LK99 considering two distinct locations for copper substitution, denoted Pb(1) and Pb(2), in the lead-phosphate apatiteGriffin (2023), Pb10(PO4)6X2, where X=O, OH. In each case, the existence of nearly-flat bands at the Fermi energy was demonstrated. The presence of such bands increases the density of states at the Fermi energy increasing the possibility of a mechanism for high-temperature superconductivity.
In the intervening months many questions about the nature of electronic states in LK99 were raised. Additional studies have been published with remarkable speed and detailCabezas-Escares et al. (2023); Hao and Fu (2024); Jiang et al. ; Lai et al. (2024); Korotin et al. (2023); Lee and Dai (2023); Shen et al. (2023), providing evidence that suggest LK99 is not a high-temperature superconductor. It is important to note that despite being one of the most common mechanisms for superconductivity, the electron-phonon coupling in LK99 has not been examined in detail for each molecular configuration. We are aware of only one work that has directly examined electron-phonon coupling strengthPaudyal et al. (2023). This is expected given the computational demands of such a quantity as well as the fact that many available density functional theory packages do not offer the capability for computation of electron-phonon coupling strength if a non-zero value of the Hubbard U is specified for a constituent atom in the compoundGiustino (2017). Many works have instead chosen to examine alternative indicators of superconductivity which are computationally less expensive such as the Fubini-Studi metricJiang et al. .
In Ref. Paudyal et al. (2023), the electron-phonon coupling strength was computed from first-principles for Pb9Cu(PO4)6O2 with Cu substituted at the Pb(2) site neglecting inclusion of a Hubbard U. It was shown that the electron-phonon coupling is weak. Even without inclusion of a Hubbard U for copper, this was a fascinating result as the electron-phonon interaction is one of the most common routes to the realization of superconductivity. It is thus a quantity of supreme importance and should be compared among the possible compositions of copper substituted lead phosphate apatite.
In this work we follow a recently established protocol to compute a proxy for the electron-phonon coupling strength, . This proxy involves computation of the electron-phonon coupling at the point only, allowing us to overcome the computational challenges that have kept these quantities absent from the literature. In Ref. Sun et al. (2022), it is shown that,
where it is reasonable to estimate as a constant for a family of materials. It was further shown in Ref. Sun et al. (2022), that the magnitude of is generally sufficient for distinguishing systems likely to support superconductivity from those in which it is highly unlikely. As a result, we have computed for Pb10(PO4)6X2 X=O, OH, with copper substitution at the Pb(1) site, allowing for a direct comparison between the compositions.
@@ II Results
Pb10(PO4)6X2 belongs to space group and for X=O, OH respectively. Details of the structure before and after copper substitution can be found in Ref. Griffin (2023). The locations for copper substitution on the two in-equivalent Pb sites, labeled Pb(1) and Pb(2) are shown in Fig. 1. The computed density of states for the relaxed structures for each of the compounds after copper substitution can be seen in Fig. 2 and the relaxed crystal structures are visible in Fig. 3. The density of states computations in Fig. 2 demonstrate that the inclusion of spin-orbit coupling has had a minimal impact below the Fermi energy, indicating that neglecting spin-orbit coupling will not invalidate the results obtained. For further information regarding the crystal structure, bulk band structure, and more, we refer the reader to Refs. Griffin (2023); Cabezas-Escares et al. (2023); Shen et al. (2023); Jiang et al. .
The resulting phonon modes at the location are shown for Pb9Cu(PO4)6X2 with Cu at the Pb(1) site for X=OH, O in Fig. 3. We immediately note that the absence of negative frequency modes for X=O, indicating that this compound is dynamically stable. In contrast, for X=OH negative phonon modes can be found. While this can indicate a dynamic instability, it is likely that such negative frequency modes would be removed through imposition of a stricter convergence threshold when relaxing the structure. This presents a computational challenge due to the lightness of the hydrogen atoms whose positions can be easily adjusted with minimal effects on the energy of the structure, slowing convergence.
The computed electron-phonon coupling strength at the location at the Fermi energy as a function of the smearing values utilized in the computation is shown in Fig. 4. In order to estimate for each compound, we utilize the value obtained in Ref. Paudyal et al. (2023) for Pb9Cu(PO4)6O2 with Cu substituted at the Pb(2) site. We then determine , where
@@
In Ref. Paudyal et al. (2023), a superconducting critical temperature of was computed for Pb9Cu(PO4)6O2 with Cu substituted at the Pb(2) site. Given the current results, a similar value should be obtained for substitution at the Pb(1) site.